OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=(x-7)/(x+3) ; g(x)=(-3x-7)/(x-1)

OpenStudy (loser66):

from f(x) , where you see x , replace by g(x) , after simplify, you will see it is = x do the same with g(f(x) , where you see x in g( ), put f(x) in, after simplify, you will see it is = x , too.

OpenStudy (loser66):

again DAT SIT

OpenStudy (shamil98):

f(x) = (x-7)(x-3) y = (x-7)/(x-3) if f and g are inverses find the inverse of f and see if they match. y = (x-7)/(x-3) x = (y-7)/(y-3) solve for y.

OpenStudy (loser66):

w.....w,,,,,,wwwut??? hehehehe... it's new to me.

OpenStudy (anonymous):

((-3x-7)/(x-1)-7)/((-3x-7)/(x-1)+3 )=x g(f(x))=(-3(x-7)/(x+3)-7)/((x-7)/(x+3)-1)=x Ok now what?

OpenStudy (shamil98):

what is that

OpenStudy (loser66):

@jessicak2 It is not DAT SIT to me, hehehe. wait for him

OpenStudy (anonymous):

I plugged in one function into the other then solved and vice versa?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm sorry, I don't understand. Yes to what?

OpenStudy (loser66):

your inverse is not right, if f (x) = \(\Large \frac{x-7}{x-3}\) its inverse g(x ) must be g(x)=\(\Large\dfrac{3x-7}{x-1}\)

OpenStudy (loser66):

Please check, because with a wrong problem, we cannot get the right answer

OpenStudy (loser66):

One more thing, Your problem is "CONFIRM" ; it implicitly indicates that f and g are inverse to each other.

OpenStudy (anonymous):

Oh my goodness I put the wrong sign!?

OpenStudy (loser66):

If it is "CONSIDER", we can solve as @shamil98 suggested

OpenStudy (loser66):

oh yea, I can say DAT SIT again, hehehehe... good luck, just do it, it 's not hard

OpenStudy (anonymous):

Thanks I got a 30/33 :)

OpenStudy (loser66):

good good good. congrats